OpenStudy (anonymous):

find the pattern for the sequence 384,192,96,48

OpenStudy (anonymous):

each consecutive term is half of the previous so the progression would be \[T _{n} = 384 \times (\frac{ 1 }{ 2 })^{n-1} \] where n is any given term is the sequence

OpenStudy (anonymous):

for example to find to third term: \[T _{3}= 384 \times (\frac{ 1 }{ 2 })^{3-1}\] \[T _{3}= 384 \times (\frac{ 1 }{ 2 })^{2}\] \[T _{3}= 384 \times \frac{ 1 }{ 4 }\] \[T _{3}= 96\]

OpenStudy (anonymous):

thank you so much!!!